### v5

# OmniRouter: Budget and Performance Controllable Multi-LLM Routing

Abstract
Large language models (LLMs) deliver superior performance but require substantial computational resources and operate with relatively low efficiency, while smaller models can efficiently handle simpler tasks with fewer resources. LLM routing is a crucial paradigm that dynamically selects the most suitable large language models from a pool of candidates to process diverse inputs, ensuring optimal resource utilization while maintaining response quality. Existing routing frameworks typically model this as a locally optimal decision-making problem, selecting the presumed best-fit LLM for each query individually, which overlook global budget constraints, resulting in ineffective resource allocation. To tackle this problem, we introduce OmniRouter, a fundamentally controllable routing framework for multi-LLM serving. Instead of making per-query greedy choices, OmniRouter models the routing task as a constrained optimization problem, assigning models that minimize total cost while ensuring the required performance level. Specifically, a hybrid retrieval-augmented predictor is designed to predict the capabilities and costs of LLMs. After obtaining the predicted cost and performance, we utilize a constrained optimizer for cost-optimal assignments that employs Lagrangian dual decomposition with adaptive multipliers. It iteratively converges toward the globally optimal query-model allocation, dynamically balancing latency minimization against quality thresholds while adhering to heterogeneous capacity constraints. Besides, we also introduce a new dataset, OmniRouteEval, which diverges from existing datasets that predominantly focus solely on either response quality or budget cost. Our dataset uniquely balances efficacy and efficiency considerations, containing sample-wise response quality and token usage measurements across 10 different LLMs, enabling comprehensive evaluation of both response quality and computational budget constraints. Experiments show that OmniRouter achieves up to 6.30% improvement in response accuracy while simultaneously reducing computational costs by at least 10.15% compared to competitive router baselines. The code and the dataset are available at https://github.com/agiresearch/OmniRouter.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities, powering a diverse range of applications from chatbots (Achiam et al., 2023; Team et al., 2023; Dubey et al., 2024; Guo et al., 2025) and code assistants (Hui et al., 2024; Wei et al., 2023; Nijkamp et al., 2023; Zhu et al., 2024) to computer-use agents. This success has spurred the widespread integration of LLMs into modern systems (Duan et al., ; Sun et al., 2024; Mei et al., 2024; Packer et al., 2023; Shi et al., 2025; Zheng et al., 2023b; Kwon et al., 2023). As LLM inference accelerates through hardware improvements, these LLM-integrated systems increasingly deploy not just one, but a set of LLMs with varying sizes, capabilities, and speeds as serving endpoints. This multi-LLM paradigm necessitates intelligent routing: the crucial task of directing incoming user queries to the most appropriate LLM instance to balance performance goals and resource efficiency.
The design of effective LLM routers has become an active area of research. Recent works have proposed various routing strategies to optimize for cost (Ong et al., 2024; Somerstep et al., 2025; Anonymous, 2024), latency (Jin et al., 2023; Stojkovic et al., 2024; Zheng et al., 2024), and performance (Zhang et al., 2025; Yue et al., 2024; Chen et al., 2024; Zhuang et al., 2024). Most existing LLM routing frameworks predominantly treat routing as a sequence of independent, greedy decisions. For each incoming query, these routers select a model based on a local optimization criterion (e.g., lowest predicted latency, cheapest model predicted to succeed) without considering system-wide resource limitations or overall performance targets. These approaches fundamentally fail to achieve Pareto-efficient resource allocation across diverse query workloads distributed among multiple LLMs. When operating under budget constraints, these greedy routers cannot globally optimize the performance-resource tradeoff, leading to suboptimal overall system efficacy. For instance, expending computational resources on marginally improving responses to simple queries may leave insufficient capacity for complex queries where performance improvements would be more valuable, which can been in Figure 1. More critically, these localized decision frameworks cannot effectively enforce global constraints such as maintaining target response quality levels while adhering to strict resource budgets, resulting in either performance degradation or resource over-utilization when deployed at scale. In summary, these approaches confront a fundamental optimization dilemma: maintaining optimal overall performance while operating under computational resource constraints and stringent budgetary limitations.
Figure 1: Comparison between traditional greedy routers and OmniRouter. Left: Greedy routers select models based on per-query optimization, leading to suboptimal allocations where Q1 (simple query) comes first and is assigned to the strong LLM, blocking Q2 (complex query) from accessing it. As a result, Q2 fails when assigned to the weak LLM. Right: OmniRouter employs constrained optimization to consider the global query distribution and model capabilities. It assigns the simple query to LLM1 (sufficient for the task) and reserves LLM2 for the complex query that requires more capability, thus maximizing overall success rate.
Recognizing these limitations, we introduce OmniRouter, a controllable routing framework that fundamentally reimagines how query-model assignments should be determined. Unlike existing approaches, OmniRouter formalizes routing in multi-LLM systems as a constrained optimization problem (Zhang et al., 2020; Bertsekas, 2014; Homaifar et al., 1994) with global performance requirements and operational constraints and proposes a two-stage routing solution. Inspired by the success of retrieval-augmented generation (RAG) in reducing hallucination of LLMs, we design a retrieval-augmented predictor at the first stage to enhance the prediction-based router. At the second stage, we design the constrained optimizer which employs a dual gradient-based approach to navigate the solution space by adjusting model selections based on quality requirements and model concurrency. Our contributions are as follows:
$\bullet$  We introduce OmniRouteEval, a grounded dataset with query-model response correctness and token usage collected across 10 different models, enabling realistic evaluation of routing strategies.
$\bullet$  We propose OmniRouter, a routing framework for multi-LLM serving that fundamentally regards routing as a constrained optimization problem rather than a series of greedy decisions, enabling system-level control of both performance and budget constraints.
$\bullet$  Extensive experiments show that OmniRouter can outperform competitive baselines across various serving scenarios and demonstrate significant improvements in response quality (up to 6.30%), cost efficiency (at least 10.15%), especially under tight operational constraints.
2 Problem Formalization
Prior methods (Ong et al., 2024; Fu et al., 2024; Parkar et al., 2024) typically score or rank models for each query independently based on heuristics such as cost-effectiveness or response likelihood. While effective in simple settings, such formulations struggle to account for system-level constraints such as limited model concurrency, global quality targets, and budget bounds that are critical in real-world multi-LLM deployments. Different from existing works that formulate LLM routing as a ranking or selection problem, our approach models it as a constrained optimization problem.
Formally, given $N$ queries and $M$ models, let $a_{i,j}\in[0,1]$ denote the capability of whether the model $j$ can successfully answer query $i$, and $c_{i,j}$ denote the money cost for answering this query. Each model $j$ is subject to a concurrency limit $L_{j}$, and the overall system must maintain a minimum average performance constraint $\alpha$ across all $N$ queries to ensure response quality. Our objective is to assign queries to models such that the total cost is minimized, while ensuring that the overall performance exceeds $\alpha$ and that no model exceeds its concurrency constraint. This problem can be formulated as the following:
$\min_{x}\quad$ $\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}\quad s.t.\quad\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha\quad\sum_{i=1}^{N}x_{i,j}\leq L_{j},\forall j\quad\sum_{j=1}^{M}x_{i,j}=1,\forall i$ (1)
It is a constrained optimization problem (Bertsekas, 2014; Homaifar et al., 1994) with global constraint $\alpha$ and local constraint $L$. We will elaborate how to solve this problem in the following section.
3 Methodology
As the capability $a_{i,j}$ and computational cost $c_{i,j}$ for each model-query pair are inherently uncertain at routing decision time. This uncertainty makes direct optimization of the routing variable $x_{i,j}$ particularly challenging. Joint optimization couples errors between prediction and allocation (Feng et al., 2023) and make the routing struggle with adaptability when query distributions shift or new models (Yu et al., 2022; Qin et al., 2023; Zhang et al., 2023). To address these challenges, we propose a two-stage approach that decouples the prediction of uncertain variables: $a_{i,j}$ and $c_{i,j}$ from the optimization of routing decisions $x_{i,j}$.
3.1 Retrieval-augmented Predictor
Inspired by retrieval-augmented generation (RAG) systems (Lewis et al., 2020; Borgeaud et al., 2022), which enhance model outputs by incorporating relevant retrieved information, we design a retrieval-augmented predictor, as illustrated in Figure 2, to integrate the generalization capabilities of trained models with the retrieval precision of historical query-model data.
We employ bert-base-uncased (Devlin, 2018) as our embedding encoder due to its representation capabilities and computational efficiency. This encoder processes both queries and LLM descriptions to generate embedding vectors $E_{q}$ and $E_{l}$ respectively, which serve as the foundation for both our training-based and retrieval-based prediction components. For the training-based component, we design a dual-head architecture built upon the embedding encoder. The first head focuses on model capability prediction, estimating a capability score $a_{i,j}^{pred}\in[0,1]$ for each query $i$ and model $j$ through:
$a_{i,j}^{pred}=\sigma(\mathbf{W}_{1}(E_{q}^{i}\cdot E_{l}^{j})+\mathbf{b}_{1})$ (2)
where $\sigma(\cdot)$ denotes the sigmoid activation function, $\mathbf{W}_{1}$ and $\mathbf{b}_{1}$ are learnable parameters. The second head performs sequence length classification by mapping the predicted output token length $l_{i,j}^{pred}$ into discrete buckets =$\mathcal{B}=\{B_{1},B_{2},...,B_{k}\}$, where $k=\lceil\frac{l_{max}}{bs}\rceil$, $l_{max}$ is the maximum sequence length, and $bs$ is the bucket size. The probability of length $l_{i,j}^{pred}$ is computed as:
$l_{i,j}^{pred}=bs\cdot\text{softmax}(\mathbf{W}_{2}(E_{q}^{i}+E_{l}^{j})+\mathbf{b}_{2})_{i}$ (3)
The bucketing strategy, a technique also employed in previous works (Jin et al., 2023; Zheng et al., 2024; Fu et al., 2024), transforms the challenging task of precise token-level prediction into a more tractable bucket-level approximation. During training, we optimize these dual objectives using mean squared error (MSE) for capability prediction and cross-entropy loss for length bucket classification.
Concurrently, our retrieval-based component leverages a vector database to identify the top-$k$ similar historical queries based on cosine similarity between query embeddings. For a given query embedding $E_{q}^{i}$, let $\mathcal{Q}_{k}(E_{q}^{i})$ be the set of top-$k$ similar queries retrieved from the database. The retrieval-based capability score $a_{i,j}^{retrieve}$ and output length score $l_{i,j}^{retrieve}$ for query $i$ and model $j$ are computed as:
$l_{i,j}^{retrieve}=\frac{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})\cdot l_{m,j}}{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})},\quad a_{i,j}^{retrieve}=\frac{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})\cdot a_{m,j}}{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})}$ (4)
where $\text{sim}(E_{q}^{i},E_{q_{m}})$ denotes the cosine similarity between the current query embedding and the retrieved query embedding, while $l_{m,j}$ and $a_{m,j}$ represent the historically observed output length and capability score for retrieved query $m$ on model $j$, respectively. To obtain the final predictions, we integrate both components through an adaptive fusion mechanism for each query-model pair:
$a_{i,j}=\gamma\cdot a_{i,j}^{pred}+(1-\gamma)\cdot a_{i,j}^{retrieve},\quad c_{i,j}=\delta\cdot tp_{j}(l_{i,j}^{pred})+(1-\delta)\cdot tp_{j}(l_{i,j}^{retrieve})$ (5)
where $\gamma$ and $\delta\in[0,1]$ are learnable parameters that control the balance between trained and retrieved predictions, and $tp(\cdot)$ is a function that maps the predicted token length to a computational cost estimate based on model $j$’s token prices. This integrated retrieval-augmented architecture ensures our predictor benefits from both generalizable patterns learned during training and specific historical performance data, resulting in more accurate predictions of the uncertain variables needed for the subsequent optimization stage.
Figure 2: Illustration of OmniRouter, including the hybrid predictor and constrained optimizer.
3.2 Constrained Optimizer
At the second stage, we solve the $x_{i,j}$ with the predicted $a_{i,j}$ and $c_{i,j}$. Inspired by optimizers (Zhang et al., 2020; Bertsekas, 2014; Wang et al., 2023), we leverage the Lagrangian dual theory and introduce Lagrangian multipliers to convert the primal problem into its Lagrangian dual problem. We only provide the crucial steps in this section, and the complete induction can found in Appendix A. By introducing Lagrangian multipliers $\lambda_{1},\lambda_{2,j},\mu_{i}$, we get the Lagrangian relaxation function of the original problem as follows:
$L(x,\lambda_{1},\lambda_{2,j},\mu_{i})=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}+\lambda_{1}\left(-\frac{\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}}{N}+\alpha\right)$
$\quad\quad\quad\quad\quad\quad\quad+\sum_{j=1}^{M}\lambda_{2,j}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right)+\sum_{i=1}^{N}\mu_{i}\left(\sum_{j=1}^{M}x_{i,j}-1\right)$ (6)
The dual function is formulated as Eq 7:
$g(\lambda_{1},\lambda_{2})=\min_{x_{i,j}}\mathcal{L}(x,\lambda_{1},\lambda_{2})=\min_{x_{i,j}}\{\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})+\lambda_{1}\alpha-\sum_{j=1}^{M}\lambda_{2,j}L_{j}\}$ (7)
and the dual problem is formulated as:
$\max_{\lambda_{1},\lambda_{2}}g(\lambda_{1},\lambda_{2})\quad\text{s.t.}\quad\lambda_{1}\geq 0,\lambda_{2,j}\geq 0,\forall j$ (8)
Then we can use gradient descent to update the partial derivative of $\lambda_{1}$ and $\lambda_{2}$, respectively, which is shown in Eq 9 and Eq 10.
$\lambda_{1}^{t+1}=\max(\lambda_{1}^{t}+\alpha_{1}(\frac{-\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}a_{i,j}}{N}+\alpha),0)$ (9)
$\lambda_{2,j}^{t+1}=\max(\lambda_{2,j}^{t}+\alpha_{2}(\sum_{i=1}^{N}x_{i,j}-L_{j}),0),\forall j\in[M]$ (10)
By updating $\lambda_{1}$ and $\lambda_{2}$, we can solve the partial sub-problem to obtain $x_{i,j}$ accordingly:
$x_{i,j}=\begin{cases}1&\text{if }j=j^{*}_{i}\\ 0&\text{otherwise}\end{cases}\quad j^{*}_{i}=\arg\min_{j\in[M]}(c_{ij}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})$ (11)
An intuitive explanation of this process is: For $\lambda_{1}$: its partial derivative represents how much the quality constraint is violated. If the average quality is less than the threshold $\alpha$, the constraint is violated, and the gradient is positive, pushing $\lambda_{1}$ to increase, which in turn encourages the scheduler to select higher-quality models in the next iteration. For $\lambda_{2,j}$, its partial derivative shows the workload violation for each model j. If a model’s assigned queries exceed its capacity, the gradient is positive, increasing $\lambda_{2,j}$, which makes this overloaded model less attractive in subsequent iterations.
Proposition 1.
Given the Lagrangian dual problem with multipliers $\lambda_{1}\geq 0$ and $\lambda_{2,j}\geq 0$ for $j\in[M]$, the optimal assignment $x^{*}_{i,j}$ for the constrained optimization problem is uniquely determined by:
$x^{*}_{i,j}=\begin{cases}1&\text{if }j=\arg\min_{j^{\prime}\in[M]}\left(c_{i,j^{\prime}}-\frac{\lambda_{1}^{*}a_{i,j^{\prime}}}{N}+\lambda_{2,j^{\prime}}^{*}\right)\\ 0&\text{otherwise}\end{cases}$ (12)
where $c_{i,j}$ and $a_{i,j}$ are the cost and capability scores predicted by the hybrid predictor, and $\lambda_{1}^{*}$ regulates the trade-off between efficiency and efficacy.
The optimal assignment rule shows that each query $i$ is routed to the model $j$ that minimizes the adjusted cost, where the adjustment term $\frac{\lambda_{1}^{*}a_{i,j}}{N}$ incentivizes selecting models with higher capability scores when $\lambda_{1}^{*}>0$.
Proposition 2.
Given the optimal dual variable $\lambda_{1}^{*}$ obtained from the gradient ascent update in Eq. (9), if $\lambda_{1}^{*}>0$, then the quality constraint is active and satisfied with equality:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}^{*}=\alpha$ (13)
where $\alpha$ is the specified quality threshold in the constrained optimization problem.
This proposition ensures that the routing algorithm achieves exactly the required quality level $\alpha$ when the quality constraint is binding, preventing over-provisioning that would unnecessarily increase costs while maintaining the desired efficacy level. Detailed proof can be referred as Appendix A.
4 Evaluation
In this section, we propose the following research questions regarding the performance of OmniRouter and conduct experiments to answer these research questions.
[…]
4.1 Datasets
Table 1: Comparison of OmniRouteEval with existing datasets used for LLM routing.
Dataset Token Usage Response Quality Cross-LLM Question Difficulty LMSYS-Chat (Zheng et al., 2023a) ShareGPT [^1] https://sharegpt.com/ RouterEval (Huang et al., 2025) OmniRouteEval
Since existing datasets for LLM routing can not satisfy the needs of routing under multiple constraints, we introduce OmniRouteEval, a dataset to facilitate the evaluation of LLM routing. The differences between OmniRouteEval and existing datasets for LLM routing are highlighted in Table 1. Within OmniRouteEval, we collect 2.7k questions sourced from established knowledge and mathematical reasoning datasets, including MMLU (Hendrycks et al., 2020), GPQA (Rein et al., 2023), MATH (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021). We select 10 different models, including 5 relatively weak models, i.e., Qwen2.5 (7B-Instruct, 14B-Instruct, 32B-Instruct) (Yang et al., 2024), Gemma2 (9B-it, 27B-it) (Team et al., 2024) and 5 relatively strong models, i.e., Qwen2.5-72B-Instruct, GPT 4o-mini, GPT-4o (Achiam et al., 2023), Gemini-1.5-flash (Team et al., 2023), Claude-3.5-sonnet  [^2] https://claude.ai/. And we collect the response correctness and token usage of these models to answer these questions using Llama-3.1-70B-Instruct (Dubey et al., 2024) as the evaluation judge. The Llama-3.1-70B is excluded from the LLM candidate pool to reduce potential biases. Further details on the OmniRouteEval dataset construction and correctness assessment methodology are provided in Appendix B.
4.2 Setup
We conduct our experiments on an Ubuntu 22.04 machine equipped with 8 RTX A5000 GPUs to serve multiple LLMs using Ollama  [^3] https://ollama.com/. During the implementation, we set the maximum output length to 1024 tokens to prevent the influence of extremely long output sequences, set top $k$ as 16 during retrieval, and $\gamma$ and $\delta$ as 0.5 during aggregation. To calculate money costs, we refer to Litellm’s costmap for price reference. The detailed price map can be found in Appendix B. Our experiments are conducted on the continuous-batching setting (Kwon et al., 2023). The serving system processes requests by dynamically forming batches based on the incoming query traffic, providing a more realistic evaluation environment for routing frameworks. In our simulation, we model real-world traffic patterns by randomly adding $n\in\{1,2,3,4\}$ queries to the queue every 0.1 seconds and performing routing decisions at 1-second intervals. If not specifically mentioned, we set the two constraints $\alpha=0.75$ and the concurrent workload constraint $L=4$, which are applied uniformly across all LLMs. Controllability analysis of different values of $\alpha$ and $L$ is presented in Section 4.4. We evaluate OmniRouter against the following routing baselines:
Cost-oriented: We adapt S3 (Jin et al., 2023) and PO (Zheng et al., 2024), which are originally designed for latency optimization as cost-oriented baselines. Specifically, S3 employs DistilBERT (Sanh et al., 2019) and PO employs Vicuna-7B (Chiang et al., 2023) as output token length predictors, respectively.
Performance-oriented: For performance-focused routing, we implement EmbedLLM (Zhuang et al., 2024) and RouterDC (Chen et al., 2024). EmbedLLM leverages an encoder-decoder architecture to embed LLM representations, while RouterDC employs contrastive learning to model query-LLM relationships, both aiming to maximize response quality regardless of computational cost.
Cost-performance Coordinated: For baselines which balance both cost and performance objectives, we employ Hybrid-LLM (Anonymous, 2024), which constructs a probabilistic router, and CARROT (Somerstep et al., 2025), which implements dual predictors using Llama3-8B (Dubey et al., 2024) as specified in its original implementation.
4.3 RQ1: Routing Performance
Table 2: Routing performance comparison, where we use accuracy, i.e., whether the routed LLM successfully answers the assigned query, and money cost of calling respective LLMs.
Metric S3 PO EmbedLLM RouterDC Hybrid-LLM CARROT OmniRouter Accuracy ($\uparrow$) 69.45% 68.71% 72.96% 73.89% 71.48% 72.41% 75.19% \$ Cost ($\downarrow$) 0.0585 0.0610 0.0896 0.0874 0.0722 0.0680 0.0515
The comparison of overall routing performance can be observed from Table 2. OmniRouter demonstrates substantial advantages over all baselines across both metrics. Given the theoretical performance bounds (lower bound is 57.41% when the worst LLM for each query is selected and upper bound is 90% when the best LLM for each query is selected), the improvement achieved by OmniRouter (1.3% to 6.48% absolute improvement) is particularly significant, which represents a meaningful advancement within this constrained optimization space. When compared to cost-oriented baselines (S3 and PO), OmniRouter achieves significantly higher success rates (5.74% and 6.48% improvements respectively) while maintaining comparable or even lower costs. Against performance-oriented approaches (EmbedLLM and RouterDC), our method achieves both higher success rates and substantially lower costs (approximately 41% cost reduction compared to RouterDC). Even when compared to cost-performance coordinated methods (Hybrid-LLM and CARROT), OmniRouter demonstrates superior performance in both dimensions.
Figure 3: Distribution of OmniRouter’s query routing decisions across difficulty levels and model capabilities.
Table 3: Performance of predictors on model capability and output token length prediction.
Method Capability Acc. Length Bucket Acc. Exact Match ±$\pm 1$ S3 0.333 0.656 PO 0.325 0.683 EmbedLLM 0.732 RouterDC 0.761 Hybrid-LLM 0.708 CARROT 0.745 0.312 0.672 OmniRouter 0.813 0.452 0.806
Performance of Predictors. We also compare the effectiveness of OmniRouter’s predictor with the baselines we have mentioned. Table 3 demonstrates OmniRouter’s significant advantages in prediction accuracy. Our approach achieves 81.3% capability prediction accuracy, outperforming the strongest baseline (RouterDC) by 5.2%. For length prediction, OmniRouter delivers 45.2% exact match accuracy and 80.6% relaxed (±$\pm 1$) accuracy, surpassing all alternatives by over 13.0%.
Routing Performance across Query Difficulty Levels. To assess the routing performance of OmniRouter deeper, we analyze its allocation patterns across different difficulty categories. Figure 3 reveals OmniRouter’s sophisticated routing strategy across query difficulty levels. For easy queries (78.4% of workload), the system successfully routes 49.2% of them to weak models. Medium-difficulty queries receive balanced treatment with equal distribution between model types (7.6% each). For hard queries, OmniRouter shows a strong preference for routing to strong models by a 4:1 ratio (5.1% vs 1.3%), demonstrating the capability of OmniRouter to match appropriate models.
4.4 RQ2: Controllability Analysis
We investigate how different routing approaches respond to varying operational constraints, i.e., performance and concurrent constraints.
(a) Cost vs. Performance Constraint
(b) Accuracy vs. Performance Constraint
Figure 4: Impact of performance constraint ($\alpha$) on cost efficiency and routing accuracy. As performance requirements increase, greedy methods exhibit unbounded cost growth while OmniRouter’s constraint optimization maintains controlled scaling.
Figure 4 reveals a critical limitation of existing greedy routing approaches: their inability to bound costs under stringent performance requirements. As the performance constraint ($\alpha$) increases from 0.70 to 0.90, baseline methods exhibit exponential cost growth: at $\alpha=0.90$, the cost of EmbedLLM reaches \$0.229, more than four times OmniRouter’s cost of \$0.054, creating a substantial cost gap ($\Delta=0.175$). This cost explosion occurs because greedy routing strategies inherently prioritize immediate performance gains when performance thresholds rise, defaulting to routing increasingly more queries to the most powerful (and expensive) models regardless of actual query complexity. In contrast, OmniRouter demonstrates remarkably controlled cost scaling, with only a 48% increase over the same range and even a slight cost reduction at $\alpha>0.85$.
(a) Cost vs. Concurrency Constraint
(b) Accuracy vs. Concurrency Constraint
Figure 5: Impact of concurrency constraint ($L$) on cost efficiency and routing accuracy. As available parallelism decreases, greedy methods struggle to make effective compromises, while OmniRouter’s constraint optimization maintains balanced allocations.
Similarly, in Figure 5, it demonstrates system behavior under varying concurrency constraints, revealing another key weakness of greedy approaches. As available concurrency decreases from 8 to 1, baseline methods show substantial cost increases (EmbedLLM: 159%, RouterDC: 147%) and significant performance degradation. When concurrency is highly limited ($L=1$), the performance gap between OmniRouter and the weakest baseline (S3) becomes dramatic ($\Delta=11.7\%$). This occurs because greedy strategies struggle to make effective compromises when resource constraints tighten. By constrast, OmniRouter substantially performs better, maintaining 73.8% accuracy even at $L=1$ while experiencing only moderate cost increases.
4.5 RQ3: Ablation Studies
Effects of Removing Modules in the Predictor. To evaluate the contribution of each component in our retrieval-augmented predictor, we conduct ablation studies by removing key modules from the full model. Results are presented in Table 4.
Table 4: Effects of removing different modules in the predictor on routing performance.
Predictor Variant Capability Acc ($\uparrow$) Bucket Acc ($\uparrow$) Accuracy ($\uparrow$) \$ Cost ($\downarrow$) Full 0.813 0.806 75.19% 0.0515 w/o Retrieval 0.751 0.724 71.85% 0.0698 w/o Training 0.728 0.682 73.33% 0.0642
Removing the retrieval component leads to a substantial drop in both capability accuracy (7.6% decrease from 0.813 to 0.751) and length prediction accuracy (10.2% decrease from 0.806 to 0.724). This degradation directly impacts routing effectiveness, resulting in lower success rate (4.4% decrease) and significantly higher operational costs (35.5% increase from \$0.0515 to \$0.0698). These findings highlight that historical query information provides critical contextual signals that enhance prediction. On the other hand, after removing the training-based component, although performance accuracy sees a larger decline (10.5% decrease to 0.728) compared to the no-retrieval variant, the success rate decreases less dramatically (only 2.5% reduction). This suggests that the retrieval mechanism alone sometimes can maintain reasonable routing decisions, which further validate our retrieval-augmented choice to enhance predictors.
Effects of Key Parameters within OmniRouter’s Predictors. We also investigate the sensitivity of OmniRouter’s routing effectiveness to key design parameters within its prediction components. However, due to the space limit, we move the results to Appendix C.
Broader Impact. By reducing the computational resources needed to achieve high-quality responses, OmniRouter contributes to more environmentally sustainable AI deployments with lower carbon footprints. While OmniRouter itself is designed to optimize resource allocation, the efficiency gains it provides might accelerate the proliferation of LLM deployments, potentially exacerbating privacy and security concerns associated with widespread AI adoption. We view these considerations not as arguments against developing more efficient routing systems, but rather as important areas for ongoing research and thoughtful implementation practices. Future work should explore how to incorporate fairness metrics directly into the constraint optimization framework.
5 Related Work
LLM Routing. To optimize performance and inference cost in multi-LLM serving systems, researchers have developed increasingly sophisticated routing approaches. RouteLLM (Ong et al., 2024) and CARROT (Somerstep et al., 2025) propose training small LLMs as routers to balance cost and performance. EmbedLLM (Zhuang et al., 2024) introduces a specialized encoder-decoder network for embedding LLM representations. HybridLLM (Anonymous, 2024) addresses data imbalance issues by proposing a probabilistic router to better represent different model capabilities across query types. C2MAB-V (Dai et al., 2024) treats routing as a contextual multi-armed bandit problem, enabling exploration-exploitation trade-offs in model selection. GraphRouter (Feng et al., 2024) leverages graph learning to jointly modeling the query-model, query-query, and model-model relationship for building routers. While these approaches have made substantial progress, they predominantly operate as greedy decision-makers, optimizing for individual queries without considering global system constraints or the implications for subsequent requests.
LLM Generation Length Prediction. Predicting LLM generation length is crucial for optimizing computational resources. Early attempts like Magnus (Cheng et al., 2024) employed random forest algorithms but achieved limited accuracy. Subsequent research has explored two main directions of prediction models: encoder-only models for classification (DynamoLLM (Stojkovic et al., 2024), S3 (Jin et al., 2023), TerriInfer (Hu et al., 2024), SSJF (Qiu et al., 2024b), and $\mu$3 (Qiu et al., 2024a)) and decoder-only models for generative prediction like Perception-only (PO) (Zheng et al., 2024). LTR (Fu et al., 2024) reformulated this as a ranking problem and utilized listwise ranking for predictor training.
6 Conclusion
In this paper, we introduce OmniRouter, a routing framework that fundamentally reframes LLM routing as a constrained optimization problem rather than a series of greedy decisions. We develop a two-stage approach with a hybrid predictor that accurately estimates model capabilities and costs, and a constrained optimizer that minimizes operational expenses while satisfying performance requirements. Our experiments with the OmniRouteEval dataset demonstrate that OmniRouter achieves up to 6.30% higher response accuracy while reducing costs by at least 10.15% compared to existing methods. OmniRouter maintains remarkable stability under varying constraints, precisely where greedy approaches fail. These results confirm OmniRouter’s effectiveness for more realistic environments requiring both performance guarantees and cost control.
Appendix
This appendix contains additional details for the paper: OmniRouter: Budget and Performance Controllable Multi-LLM Routing. The appendix is organized as follows: Section §A is about complete induction and proof of the constrained optimization. Section §B is about dataset details. Section §C is about additional experimental results. Section §D discusses more related works.
Appendix A Complete Induction and Proof of the Constrained Optimization
A.1 Complete Induction
The primal problem is as below:
$\min_{x}\quad$ $\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}$ (14)
s.t. $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha$ (15)
$\sum_{i=1}^{N}x_{i,j}\leq L_{j},\forall j\in[M]$ (16)
$\sum_{j=1}^{M}x_{i,j}=1,\forall i\in[N].$ (17)
we introduce three types of Lagrange multipliers as:
  • $\lambda_{1}\geq 0$ for the quality constraint (inequality)
  • $\lambda_{2,j}\geq 0$ for each capacity constraint (inequality)
  • $\mu_{i}$ for each assignment constraint (equality)
and write the Lagrangian function as:
$L(x,\lambda_{1},\lambda_{2,j},\mu_{i})$ (18)
$=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}+\lambda_{1}\left(-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}+\alpha\right)$ (19)
$\quad+\sum_{j=1}^{M}\lambda_{2,j}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right)+\sum_{i=1}^{N}\mu_{i}\left(\sum_{j=1}^{M}x_{i,j}-1\right)$ (20)
which can rearranged to group terms with $x_{i,j}$:
$L(x,\lambda_{1},\lambda_{2,j},\mu_{i})=\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}\left(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}+\mu_{i}\right)+\lambda_{1}\alpha-\sum_{j=1}^{M}\lambda_{2,j}L_{j}-\sum_{i=1}^{N}\mu_{i}$ (21)
The KKT optimality conditions for this problem are: Stationarity:
$\frac{\partial L}{\partial x_{i,j}}=c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}+\mu_{i}=0,\quad\forall i,j$ (22)
Primal Feasibility:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha,\quad\sum_{i=1}^{N}x_{i,j}\leq L_{j},\forall j,\quad\sum_{j=1}^{M}x_{i,j}=1,\forall i$ (23)
Dual Feasibility:
$\lambda_{1}\geq 0,\quad\lambda_{2,j}\geq 0,\forall j$ (24)
Complementary Slackness:
$\lambda_{1}\left(\alpha-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\right)=0,\quad\lambda_{2,j}\left(L_{j}-\sum_{i=1}^{N}x_{i,j}\right)=0,\quad\forall j$ (25)
Note that for each $i$, we have the equality constraint $\sum_{j=1}^{M}x_{i,j}=1$. This implies that for each query $i$, exactly one LLM $j$ must be selected. From the stationarity condition, for any fixed $i$, comparing two different indices $j$ and $k$:
$c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}+\mu_{i}=0$ $c_{i,k}-\frac{\lambda_{1}a_{i,k}}{N}+\lambda_{2,k}+\mu_{i}=0$ (26)
Subtracting these equations eliminates $\mu_{i}$:
$(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})=(c_{i,k}-\frac{\lambda_{1}a_{i,k}}{N}+\lambda_{2,k})$ (27)
This implies that for a given $i$, the optimal solution should choose the $j$ that minimizes $(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})$.
Then the dual function becomes:
$g(\lambda_{1},\lambda_{2})=\min_{x_{i,j}}\{\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})+\lambda_{1}\alpha-\sum_{j=1}^{M}\lambda_{2,j}L_{j}\}$ (28)
Note that $\mu_{i}$ has disappeared from the dual function because we’ve analytically incorporated the equality constraints. The dual problem can now be written as:
[…]
The partial derivatives for the remaining multipliers are: For $\lambda_{1}$:
$\frac{\partial L}{\partial\lambda_{1}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}a_{i,j}+\alpha$ (30)
For $\lambda_{2,j}$:
$\frac{\partial L}{\partial\lambda_{2,j}}=\sum_{i=1}^{N}x_{i,j}-L_{j},\quad\forall j\in[M]$ (31)
The gradient ascent update rules are:
$\lambda_{1}^{t+1}=$ $\max\left(\lambda_{1}^{t}+\alpha_{1}\left(-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}a_{i,j}+\alpha\right),0\right)$ (32)
$\lambda_{2,j}^{t+1}=$ $\max\left(\lambda_{2,j}^{t}+\alpha_{2}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right),0\right),\quad\forall j\in[M]$ (33)
For fixed multipliers, the optimal assignment for each $i$ is:
$x_{i,j}=\begin{cases}1&\text{if }j=j^{*}_{i}\\ 0&\text{otherwise}\end{cases}$ (34)
where
$j^{*}_{i}=\arg\min_{j\in[M]}\left(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}\right)$ (35)
  • $\lambda_{1}$ acts as a penalty for violating the quality constraint. When the average quality is below $\alpha$, $\lambda_{1}$ increases, encouraging selection of higher-quality options.
  • $\lambda_{2,j}$ penalizes capacity violations for each model $j$. When a model exceeds its capacity $L_{j}$, its corresponding multiplier increases, making it less attractive in subsequent iterations.
  • The equality constraints ($\mu_{i}$) are handled analytically by direct substitution, which simplifies the dual problem.
  • The algorithm alternates between updating multipliers and assignments until convergence.
A.2 Proof
Proposition 1. Given the Lagrangian dual problem with multipliers $\lambda_{1}\geq 0$ and $\lambda_{2,j}\geq 0$ for $j\in[M]$, the optimal assignment $x^{*}_{i,j}$ for the constrained optimization problem is uniquely determined by:
$x^{*}_{i,j}=\begin{cases}1&\text{if }j=\arg\min_{j^{\prime}\in[M]}\left(c_{i,j^{\prime}}-\frac{\lambda_{1}^{*}a_{i,j^{\prime}}}{N}+\lambda_{2,j^{\prime}}^{*}\right)\\ 0&\text{otherwise}\end{cases}$ (36)
Proof.
We start with the original optimization problem:
$\min_{x}\quad$ $\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}$ (37)
s.t. $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha$ (38)
$\sum_{i=1}^{N}x_{i,j}\leq L_{j},\quad\forall j\in[M]$ (39)
$\sum_{j=1}^{M}x_{i,j}=1,\quad\forall i\in[N]$ (40)
$x_{i,j}\in\{0,1\},\quad\forall i,j$ (41)
The Lagrangian function is:
$L(x,\lambda_{1},\lambda_{2},\mu)$ $=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}+\lambda_{1}\left(\alpha-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\right)$ (42)
$\quad+\sum_{j=1}^{M}\lambda_{2,j}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right)+\sum_{i=1}^{N}\mu_{i}\left(\sum_{j=1}^{M}x_{i,j}-1\right)$ (43)
From the KKT stationarity condition:
$\frac{\partial L}{\partial x_{i,j}}=c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}+\mu_{i}=0$ (44)
where $\mu_{i}$ is the Lagrangian multiplier for the constraint $\sum_{j=1}^{M}x_{i,j}=1$.
Rearranging the stationarity condition:
$\mu_{i}=-\left(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}\right)$ (45)
Since $\sum_{j=1}^{M}x_{i,j}=1$ and $x_{i,j}\in\{0,1\}$, exactly one $x_{i,j}$ equals 1 for each query $i$. For the optimal solution, we choose $j^{*}$ that minimizes the expression in the stationarity condition:
$j^{*}=\arg\min_{j\in[M]}\left(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}\right)$ (46)
Therefore, $x_{i,j^{*}}=1$ and $x_{i,j}=0$ for $j\neq j^{*}$, which completes the proof. ∎
Proposition 2. Given the optimal dual variable $\lambda_{1}^{*}$ obtained from the gradient ascent update, if $\lambda_{1}^{*}>0$, then the quality constraint is active and satisfied with equality:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}^{*}=\alpha$ (47)
Proof.
From the complementary slackness condition in the KKT conditions:
$\lambda_{1}^{*}\left(\alpha-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}^{*}\right)=0$ (48)
This equation states that either:
  • (i)
    $\lambda_{1}^{*}=0$, or
  • (ii)
    $\alpha-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}^{*}=0$
Given that $\lambda_{1}^{*}>0$, condition (i) is false, thus condition (ii) must hold:
$\alpha-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}^{*}=0$ (49)
Rearranging yields:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}^{*}=\alpha$ (50)
This demonstrates that when $\lambda_{1}^{*}>0$, the quality constraint is satisfied with equality. The dual variable $\lambda_{1}^{*}$ acts as a ”price” for quality: the optimizer finds the solution that exactly meets the quality threshold $\alpha$ while minimizing cost. ∎
Appendix B Dataset Details
B.1 Statistics of the Dataset
The statistics of the OmniRouteEval are in Table 5.
We also set the difficulty for each question, which is determined by the number of models that can answer the question correctly, i.e., Easy: $\{8,9,10\}$ models can answer correctly, Medium: $\{4,5,6,7\}$ models can answer correctly, Hard: $\{0,1,2,3\}$ models can answer correctly. And the proposition is: 78.4%, 15.2%, 6.4%, respectively.
Table 5: Statistics of the data distribution in our dataset.
Data Source # Samples Ratio
MMLU 1000 37.06%
GPQA-Diamond 198 7.33%
Math-500 500 18.53%
GSM8K 1000 37.06%
B.2 Cost Price Map
For calculating costs for LLMs to answer queries, we refer to litellm [^4] https://docs.litellm.ai/docs/ for calculating money cost, the money price regarding the model we use are shown in Table 6.
Table 6: Money Cost Map of LLMs
Model Name Cost per 1M Input Tokens (\$) Cost per 1M Output Tokens (\$) Qwen-2.5-7B-Instruct (Yang et al., 2024) 0.267 0.267 Qwen-2.5-14B-Instruct (Yang et al., 2024) 0.534 0.534 Qwen-2.5-32B-Instruct (Yang et al., 2024) 1.22 1.22 Qwen-2.5-72B-Instruct (Yang et al., 2024) 2.745 2.745 Gemma-2-9B-it (Team et al., 2024) 0.343 0.343 Gemma-2-27B-it (Team et al., 2024) 1.03 1.03 gpt-4o-mini (Achiam et al., 2023) 0.15 0.6 gpt-4o (Achiam et al., 2023) 2.5 10 gemini-1.5-flash (Team et al., 2023) 0.075 0.3 claude-3-5-sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet 3 15
B.3 Response Correctness Evaluation
We use the Llama-3.1-70B-Instruct (Dubey et al., 2024) as the LLM judge to evaluate whether a generated response meets the ground truth answers. The prompt is shown as below.
Prompt for using LLM as the judge. Prompt: The ground truth answer is: {gt_answer}. The prediction answer is: extracted_answer. Judge whether the prediction answer is correct or not. You just need to output ’True’ or ’False’.
Appendix C Additional Experiments
C.1 Effects of Key Parameters within OmniRouter’s Predictors.
We investigate the sensitivity of OmniRouter’s routing effectiveness to key design parameters within its prediction components. As shown in Table 7, the number of buckets used for discretizing output length predictions significantly impacts routing performance. A smaller number of buckets (10) yields substantially higher prediction accuracy (45.2% exact match, 80.6% ±$\pm 1$ accuracy), which directly leads to higher routing performance (75.19% success rate) and lower operational costs (\$0.0515). This finding suggests that coarser-grained discretization may be sufficient for effective routing decisions, while excessive granularity can introduce noise. Regarding the retrieval component, Table 8 demonstrates how varying $K$ (the number of historical samples) affects prediction accuracy and routing performance. The system achieves optimal performance at $K=16$, with the highest success rate (75.19%) and lowest cost (\$0.0515). We observe that prediction accuracy remains relatively stable across moderate $K$ values (8-32), with only slight degradation at the extremes. Too few samples ($K=4$) provide insufficient historical context, while too many ($K=64$) likely introduce noise from less relevant queries, both resulting in suboptimal routing decisions. This indicates the retrieval-augmentation requires careful calibration to balance between relevant historical information and potentially misleading outliers.
Appendix D More Related Works
Large Language Models (LLMs) have demonstrated remarkable capabilities across various domains, including mathematical reasoning (Wei et al., 2022; Yao et al., 2024; Besta et al., 2024; Jin et al., 2024), code generation (Wei et al., 2023; Nijkamp et al., 2023; Hui et al., 2024), and recommender systems  Xu et al. (2024a; b; 2023); Mei & Zhang (2023). Many studies (Wei et al., 2022; Yao et al., 2024; Besta et al., 2024; Jin et al., 2024; 2025) have demonstrated that different prompts can have great influences on the model behavior, thus making the prediction of LLM generation length challenging. Early attempts like Magnus (Cheng et al., 2024) employed random forest algorithms but achieved limited accuracy. Subsequent research has explored two main directions of prediction models: encoder-only models for classification (DynamoLLM (Stojkovic et al., 2024), S3 (Jin et al., 2023), TerriInfer (Hu et al., 2024), SSJF (Qiu et al., 2024b), and $\mu$3 (Qiu et al., 2024a)) and decoder-only models for generative prediction like Perception-only (PO) (Zheng et al., 2024). Fu et al. (Fu et al., 2024) reformulated this as a ranking problem and utilized listwise ranking for predictor training. Due to the inherent difficulty in precise output length prediction, several works (Jin et al., 2023; Zheng et al., 2024; Fu et al., 2024) adopted a bucketing approach for approximate estimation. However, these existing studies primarily emphasize computational efficiency while overlooking a critical aspect: model capability - specifically, whether a model has the capability to answer a given query. Therefore, considering both response quality and response cost to construct a general scheduling framework remains a critical challenge.
Table 7: Effects of number of buckets used in predictors on routing performance.
# Bucket Bucket Acc ($\uparrow$) Capability Acc. ($\uparrow$) \$ Cost ($\downarrow$) Exact Match ±$\pm 1$ 10 0.452 0.806 75.19% 0.0515 20 0.269 0.526 73.22% 0.0725 50 0.162 0.299 73.22% 0.834 100 0.124 0.189 70.74% 0.0962
Table 8: Effects of variations of K in historical samples on routing performance.
K Capability Acc. ($\uparrow$) Length Acc ($\uparrow$) SR ($\uparrow$) \$ Cost ($\downarrow$) Exact Match ±$\pm 1$ 4 0.773 0.430 0.784 74.07% 0.0614 8 0.785 0.443 0.786 74.81% 0.0565 16 0.806 0.452 0.806 75.19% 0.0515 32 0.785 0.440 0.776 74.81% 0.0565 64 0.780 0.422 0.764 73.70% 0.0684